"Word Problems Involving Products and Factoring"
Find two consecutive positive numbers such that the product of the sum and difference of the numbers plus 8 is the sum of their squares. can you solve this pls=( tyty so much

Answer by ankor@dixie-net.com(22740)

Find two consecutive positive numbers
x, (x+1)
:
such that the product of the sum and difference of the numbers plus 8 is the sum of their squares.
(x + (x+1))*((x+1)-x) + 8 = x^2 + (x+1)^2
:
((2x+1) * 1) + 8 = x^2 + (x^2 + 2x + 1)
:
2x + 9 = 2x^2 + 2x + 1
:
9 - 1 = 2x^2 + 2x - 2x
:
8 = 2x^2
:
4 = x^2
:
x = 2 and 3 are the numbers
:
Check in statement
"the product of the sum and difference of the numbers plus 8 is the sum of their squares.
(2+3) * (3-2) + 8 = 2^2 + 3^2
5 + 8 = 4 + 9